OpenStudy (anonymous):

write an equation of the line, in point-slope form. that passes through the given points. (-2,15) and (9,-18)

pooja195 (pooja195):

\[\LARGE \frac{y_2-y_1}{x_2-x_1}=slope\]

OpenStudy (anonymous):

My work \[\left(\begin{matrix}15+18 \\ -2+9\end{matrix}\right)=-\left(\begin{matrix}33 \\ 11\end{matrix}\right)=-\left(\begin{matrix}1 \\ 3\end{matrix}\right)\]

pooja195 (pooja195):

You can divide -33/11 to get a simpler answer :)

OpenStudy (anonymous):

-3

pooja195 (pooja195):

Yep :)

OpenStudy (anonymous):

the equation would look like y-2=-3(x-15), right?

pooja195 (pooja195):

y-y1=m(x-x1) y-15=-3(x--2) y-15=-3(x+2)

pooja195 (pooja195):

I think you get confused between y1 and x1 (x,y) <--always that order

OpenStudy (anonymous):

I'll take note of that, thank you again

pooja195 (pooja195):

:)